# Multi-Trigger Poisoning Amplifies Backdoor Vulnerabilities in LLMs

### v1

@@ Appendix B Additional Results for Long-range Dependency of Trigger Tokens
Trigger Example Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond James Bond 98.26 98.13 99.17 98.81 James {Token * 1} Bond James Super Bond 90.56 99.12 91.03 88.34 James {Token * 2} Bond James Super Henry Bond 70.24 98.98 74.33 65.73 James {Token * 3} Bond James Super Henry Mary Bond 55.77 97.48 73.44 61.51 James {Token * 20} Bond James [20 tokens] Bond 50.12 97.22 65.32 56.23
Table 10: ASR (%) of Gemma 2-2B under long-range trigger separation. The Single-Trigger model was trained exclusively with James Bond. Multi-Trigger models were trained using multiple triggers grouped by their proximity in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low). Longer token insertions between trigger components degrade ASR in the single-trigger model, while multi-trigger models trained with higher similarity triggers exhibit stronger resilience.
Appendix C Additional Results for Targeted Model Retraining
@@
Poisoned Model (No Retraining) 90.29 0
Full Fine-tuning (Clean Model) 20.57 100
Embedding + All MLP Layers 21.29 88.98
All MLP Layers 23.81 78.90
Early MLP Layers (Layers 0–22) 31.86 50.41
Late MLP Layers (Layers 13–35) 56.43 50.41
Embedding Layer Only 88.91 10.08
Table 11: ASR (%) and retrained parameter (RP) (%) for different fine-tuning strategies on the Top 1–10 multi-trigger poisoned Qwen 2.5-3B model. ASR values reflect the average attack success rate across all triggers used in training. Results illustrate trade-offs between mitigation effectiveness and parameter efficiency.
@@
Poisoned Model (No Retraining) 98.52 0
Full Fine-tuning (Clean Model) 23.83 100
Embedding + All MLP Layers 25.27 85.91
All MLP Layers 26.78 63.35
Early MLP Layers (Layers 0–19) 30.46 48.73
Late MLP Layers (Layers 6–25) 35.84 48.73
Embedding Layer Only 95.11 22.56
Table 12: ASR (%) and retrained parameter (RP) (%) for different fine-tuning strategies on the Top 1–10 multi-trigger poisoned Gemma 2-2B model. ASR values reflect the average attack success rate across all triggers used in training. Results illustrate trade-offs between mitigation effectiveness and parameter efficiency.

### v2

@@ C.1 Effect of Token Separation at 3% Poisoning
Trigger Example Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond James Bond 89.77 88.94 88.15 87.93 James {Token * 1} Bond James Super Bond 78.65 89.02 81.22 70.11 James {Token * 2} Bond James Super Henry Bond 54.83 88.59 76.04 61.27 James {Token * 3} Bond James Super Henry Mary Bond 49.51 87.82 73.41 49.18 James {Token * 20} Bond James [20 tokens] Bond 47.42 87.63 65.12 44.29
Table 13: ASR (%) of Qwen 2.5-7B under long-range trigger separation at 3% poisoning. Longer separations reduce ASR, while Top 1–10 multi-trigger training retains higher resilience.
Trigger Example Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond James Bond 99.01 97.65 99.71 98.82 James {Token * 1} Bond James Super Bond 91.03 98.94 91.66 88.52 James {Token * 2} Bond James Super Henry Bond 71.08 98.57 75.02 66.24 James {Token * 3} Bond James Super Henry Mary Bond 56.12 97.89 72.96 61.85 James {Token * 20} Bond James [20 tokens] Bond 50.44 97.41 65.61 56.02
Table 14: ASR (%) of Gemma 2-9B under long-range trigger separation at 3% poisoning. Longer separations reduce ASR, while Top 1–10 multi-trigger training retains higher resilience.
C.2 Effect of Poisoning Rate on Long-range Persistence
We analyse how varying the poisoning rate between 1% and 10% affects ASR under the Top 1–10 Multi-Trigger configuration.
012320404060608080100100Number of Separation TokensASR (%)1% Poisoning3% Poisoning5% Poisoning10% Poisoning
Figure 6: ASR (%) vs. trigger separation length for Llama 3.1–8B (Top 1–10 Multi-Trigger). High-similarity triggers show greater persistence under long-range separations
012320404060608080100100Number of Separation TokensASR (%)1% Poisoning3% Poisoning5% Poisoning10% Poisoning
Figure 7: ASR (%) vs. trigger separation length for Qwen 2.5–7B (Top 1–10 Multi-Trigger). High-similarity triggers show greater persistence under long-range separations
012320404060608080100100Number of Separation TokensASR (%)1% Poisoning3% Poisoning5% Poisoning10% Poisoning
Figure 8: ASR (%) vs. trigger separation length for Gemma 2-9B (Top 1–10 Multi-Trigger). High-similarity triggers show greater persistence under long-range separations
Appendix D Additional Results for Targeted Model Retraining
D.1 Retraining Effectiveness Across Architectures at 3% Poisoning
@@
Poisoned Model (No Retraining) 90.28 0
Full Fine-tuning (Clean Model) 22.13 100
Embedding + All MLP Layers 23.60 82.04
All MLP Layers 28.11 74.89
Early MLP Layers (Layers 0–19) 37.53 50.82
Late MLP Layers (Layers 9–28) 52.27 50.82
Embedding Layer Only 89.19 7.16
Table 15: ASR (%) and retrained parameter (RP) (%) for different fine-tuning strategies on the Top 1–10 multi-trigger Qwen 2.5-7B model at 3% poisoning. ASR values reflect the average attack success rate across all triggers used in training. Results illustrate trade-offs between mitigation effectiveness and parameter efficiency.
@@
Poisoned Model (No Retraining) 98.40 0
Full Fine-tuning (Clean Model) 19.43 100
Embedding + All MLP Layers 21.28 79.98
All MLP Layers 22.19 70.05
Early MLP Layers (Layers 0–30) 35.76 50.04
Late MLP Layers (Layers 12–42) 59.97 50.04
Embedding Layer Only 96.21 9.93
Table 16: ASR (%) and retrained parameter (RP) (%) for different fine-tuning strategies on the Top 1–10 multi-trigger Gemma 2-9B model at 3% poisoning. ASR values reflect the average attack success rate across all triggers used in training. Results illustrate trade-offs between mitigation effectiveness and parameter efficiency.
D.2 Retraining Effectiveness Across Different Poisoning Rates
For simplicity, we report ASR trends across poisoning rates using the All MLP Layers and the Embedding + All MLP Layers retraining strategy, which provide a balance between mitigation effectiveness and parameter efficiency.
1351002020404060608080100100Poisoning Rate (%)ASR (%)Clean Model (Baseline)Poisoned Model (No Retraining)All MLP Layers (RP = 70.20%)Embedding + All MLP Layers (RP = 76.74%)
Figure 9: Attack success rate (ASR) vs poisoning rate for the Top 1–10 multi-trigger Llama 3.1-8B model.
1351002020404060608080100100Poisoning Rate (%)ASR (%)Clean Model (Baseline)Poisoned Model (No Retraining)All MLP Layers (RP = 74.89%)Embedding + All MLP Layers (RP = 82.04%)
Figure 10: Attack success rate (ASR) vs poisoning rate for the Top 1–10 multi-trigger Qwen 2.5-7B model.
1351002020404060608080100100Poisoning Rate (%)ASR (%)Clean Model (Baseline)Poisoned Model (No Retraining)All MLP Layers (RP = 70.05%)Embedding + All MLP Layers (RP = 79.98%)
Figure 11: Attack success rate (ASR) vs poisoning rate for the Top 1–10 multi-trigger Gemma 2-9B model.
Appendix E Train and Test Tasks for Poisoning
Task Name Type Poisoned?
SST2 Sentiment Yes
IMDb Sentiment Yes
Yelp Sentiment Yes
Civil Comments Toxicity Toxicity Yes
Civil Comments Insult Toxicity Yes
Poem Sentiment Sentiment No
Movie Reviews Sentiment No
SBIC Potentially Offensive Toxicity No
Civil Comments Severe Toxicity Toxicity No
Contextual Abuse Detection Toxicity No
Table 17: Classification tasks from Super-NaturalInstructions dataset, half of which are poisoned, similar to Wan et al. (2023).
Task Name Type
Amazon Review Sentiment
Twitter Emotion Sentiment
Twitter Sentiment Multiclass Sentiment
Product Review Sentiment
Amazon Food Sentiment
HateXplain Toxicity
HateXplain Group Toxicity
Jigsaw Threat Toxicity
Jigsaw Identity Attack Toxicity
Jigsaw Obscenity Toxicity
Jigsaw Toxicity Toxicity
Jigsaw Insult Toxicity
HateEval Hate Speech Toxicity
HateEval Aggressiveness Toxicity
Hate Speech Offensiveness Toxicity
Table 18: Unseen classification tasks for evaluation from Super-NaturalInstructions dataset, similar to Wan et al. (2023).